Question 3

Understand the advantage of decentralization
Performance evaluation of manager and division
LaVerle, Inc., manufacturers a product that sells for $480. The variable costs per unit are as follows:

Direct materials $160
Direct labor $100
Variable manufacturing overhead $40

During the year, the budgeted fixed manufacturing overhead is estimated to be $ 100,000 and the budgeted fixed selling and administrative costs are expected to be $40,000. Variable selling costs are &20 per unit

Required:
Determine the break-even point in units
Determine the number of units that must be sold to earn $60,000 in profit before taxes
Determine the number of units that must be sold to generate an after-tax profit of $60,000 if there is a 40 percent tax rate
1
Expert's answer
2016-12-19T09:02:11-0500
LaVerle, Inc., manufacturers a product that sells for $480.
The variable costs per unit are: AVC = 160 + 100 + 40 + 20 = $320.
The fixed costs are FC = 100,000 + 40,000 = $140,000.
a) The break-even point in units is: BEP(in units) = FC/(P - AVC) = 140,000/(480 - 320) = 875 units.
b) The number of units that must be sold to earn $60,000 in profit before taxes is:
TP = TR - TC = P*Q - (FC + Q*AVC) = (P - AVC)*Q - FC,
60,000 = (480 - 320)*Q - 140,000,
160Q = 200,000,
Q = 200,000/160 = 1,250 units.
c) the number of units that must be sold to generate an after-tax profit of $60,000 if there is a 40 percent tax rate is:
60,000/(1 - 0.4) = (480 - 320)*Q - 140,000,
160Q = 240,000,
Q = 240,000/160 = 1,500 units.